# Spatial super-resolution in nanosensing with blinking emitters

### v1

@@ Abstract
We propose a method of spatial resolution enhancement in metrology (thermometry, magnetometry, pH estimation and similar methods) with blinking fluorescent nanosensors by combining sensing with super-resolution optical fluctuation imaging (SOFI). To demonstrate efficiency of this approach, a model experiment with laser diodes modeling fluctuating nanoemitters and intentional blurring of the image is performed. The 2nd, 3rd, and 4th order cumulant images provide improvement of the contrast and enable successful reconstruction of smaller features of the modeled temperature (or any other physical parameter) distribution relatively to the intensity-based approach. We believe that blinking fluorescent sensing agents being complemented with the developed image analysis technique could be utilized routinely in the life science sector for recognizing the local changes in the spectral response of blinking fluorophores, e.g. delivered targetly to the wanted cell or even organelle. It is extremely useful for the local measurements of living cells’ physical parameters changes due to applying any external “forces”, including disease effect, aging, healing or response to the treatment.
@@ I Introduction
Fluorescent nanosensors represent a powerful tool for measurement of temperature, magnetic field, pH, and other parameters with high accuracy at small scales [1, 2, 3, 4, 5, 6]. In particular, the dependence of their fluorescent signal on the mentioned parameters can be utilized for investigation of processes in living cells [2, 7, 8]. To ensure high spatial resolution, one needs to localize the nanoemitters precisely and to distinguish signals from closely located nanosensors. This requirement can be fulfilled, for example, by placement of fluorescent nanodiomonds or quantum dots at a manipulator needle [9, 10, 11, 12, 13]. This approach requires precise mechanical control and does not allow to analyze environment parameters at several points of the sample simultaneously. For a large number of nanosensors, simultaneously attached to the sample, one can apply deterministic super-resolution approaches based on selective (localized) excitation of emitters (stimulated emission depletion (STED), ground state depletion (GSD), reversible saturable optical fluorescence transitions between spin levels (spin-RESOLFT), structured illumination imaging) [14, 15, 16, 17, 18], as well as stochastic methods (stochastic optical reconstruction microscopy (STORM), photoactivated localization microscopy (PALM), points accumulation for imaging in nanoscale topography (PAINT)) [19, 20, 21, 22], ensuring independent localization of single emitters due to their blinking. Those approaches are also quite sophisticated and require the nanosensors to remain in the inactive state most of the time, thus increasing needed exposition time.
@@
The idea of combined usage of blinking statistics of fluorophores for super-resolution and certain other degree of freedom for getting additional information is somewhat close to so called “color SOFI” [29, 30, 31, 32] — super-resolution technique, which applies SOFI to several (typically two) spectral channels. However, the tasks, solved by those two approaches, are different. Color SOFI deals with discrimination between several (two or three) discrete types of fluorescent emitters, serving as biological labels. In contrast, the method proposed in this paper, deals with continuous-variable sensing: the emitters are of the same type, but their fluorescent signal differs due to variations of local environment.
@@
To demonstrate principles of the proposed approach and show its feasibility, we performed a model experiment mimicking blinking fluorescent nanoemitters. We used classical photodiods with the controlled signal imitating random switching between “bright” and “dark” states. The image was intentionally blurred by inaccurate focusing on the detection plane so that the spots produced by neighbor “emitters” were strongly overlapping. Model “temperature” of the “emitters” was encoded in the “fluorescence lifetime” — the delay between the “excitation pulse” and the emission response. The subsequent application of the proposed SOFI-based sensing approach allowed to enhance spatial resolution approximately twice.
@@
The paper is organized as follows. In Section II, we discuss basic ideas of the approach and list the procedures constituting the proposed technique of spatially super-resolving sensing. Section III describes the model experiment, data acquisition, and reconstruction results for the model “temperature” distribution. Discussion of the results and possible directions of further research are presented in Section IV.
@@ II.1 Super-resolution optical fluctuations imaging
Strictly speaking, Eqs. (2) and (4) are valid for classical signals — in the limit of a large number of detected photons, when the shot noise is small relatively to the fluctuations caused by emitters’ blinking. Detailed theoretical discussion of that assumption will be presented in Ref. [33].
To increase spatial resolution, SOFI deals with stochastic fluctuations of the total fluorescence signal, integrated over each frame duration and certain spectral range, rather than on spectral or temporal intensity distribution inside a frame. The technique itself does not impose any specific constraints on time or spectral windows, used during signal acquisition within frames. That property is crucial for the proposed sensing approach and was efficiently used in the color SOFI [29, 32].
@@ II.2 Sensing with fluorescent nanoparticles
Sensing capabilities of fluorescent nanoparticles can be provided by dependence of their fluorescence lifetime, spectrum, or total intensity on the local parameters of the environment: temperature, pH, pressure, etc. [1, 2, 34, 35, 36, 37, 38, 39] (Fig. 1). Since blinking of the emitters introduces significant noise to the total intensity of fluorescence, we focus on the cases when the environment influences distribution of the fluorescent signal over certain parameter (variable) $x$ within the frame. Let $f(\mathbf{r},t,x)$ be the density of the signal over the variable $x$. Then, the total fluorescent signal for the frame $t$ is defined as
@@
Figure 1: Sensing an environment parameter $\theta$ by its influence on fluorescence signal: changes of the spectrum (a), fluorescence lifetime (b), or total intensity (c). The latter case is incompatible with SOFI because of blinking-induced noise of absolute intensity. Nonlinear dependence of the estimator $\bar{\theta}=\Theta(\theta)$, defined by Eq. (15), can be taken into account by appropriate calibration (d). Dotted arrows schematically illustrate reconstruction of the environment parameter $\theta$ from the estimator value $\bar{\theta}$ calculated from the fluorescence signal.
@@
where the dependence on the $k$-th emitter intensity $\varepsilon_{k}(x,\theta)$ on the variable $x$ (wavelength or delay of the fluorescence signal) is affected by the local value of the environment parameter $\theta=\theta_{k}$ in the vicinity of the $k$-th emitter. To infer the local values $\theta(\mathbf{r})$ of the parameter in question, one needs to specify certain model for the dependence of $\varepsilon_{k}(x,\theta)$ on $\theta$ and, typically, to assume that all the emitters are described by the same model $\varepsilon_{k}(x,\theta_{k})=\varepsilon_{k}\varepsilon(x,\theta_{k})$ up to a constant emitter-dependent factor $\varepsilon_{k}$. To eliminate additional noise, caused by blinking of the emitters, it is worth normalizing the signal by dividing it by its integral over $x$.
First, let us assume that the dependence of the emitter’s signal on the parameter of environment is linear in the vicinity of the reference value $\theta_{0}$:
$\varepsilon(x,\theta)=S_{0}(x)+S_{1}(x)(\theta-\theta_{0}).$ (11)
As it is shown in Appendix A, one can build an estimator
@@
and ensure, by appropriate choice of the weights $Q_{0}(x)$ and $Q_{1}(x)$, that the estimator retrieves all the information contained in the intensity profile $\varepsilon(x,\theta)$ after its normalization (i.e., represents sufficient statistics).
Eq. (12) indicates the fact that the profile $\varepsilon(x,\theta)$ can be “compressed” to just two scalar characteristics (linear functionals)
[…]
which are sufficient for retrieving all the available information about the investigated parameter $\theta$. One can move further and consider the two scalars $Y_{0}$ and $Y_{1}$ for some arbitrary weights $Q_{0}(x)$ and $Q_{1}(x)$ without applying the constraints (13). Practically, those two quantities may correspond to the signal accumulated over two spectral or temporal windows. In that case (see Appendix B), one can still find two linear combinations $\bar{Y}_{0}$ and $\bar{Y}_{1}$ of the scalars so that the corresponding weights $\bar{Q}_{0}(x)$ and $\bar{Q}_{1}(x)$ satisfy Eq. (13). However, such measurement-induced choice of the weights is, generally, non-optimal and may lead to loss of information relatively to the full profile analysis. The estimator for the environment parameter $\theta$ can be constructed as
$\bar{\theta}=\theta_{0}+\frac{Y_{1}A_{00}-Y_{0}A_{10}}{Y_{0}A_{11}-Y_{1}A_{01}},$ (15)
where
[…]
represent coefficients in decomposition of $Y_{i}$ over powers of the deviation $\theta-\theta_{0}$:
$Y_{i}=A_{i0}+A_{i1}(\theta-\theta_{0}).$ (17)
In general case, the dependence of $\varepsilon(x,\theta)$ on $\theta$ is nonlinear. Considering small deviations $|\theta-\theta_{0}|$, one can linearize $\varepsilon(x,\theta)$ and define
$S_{0}(x)=\varepsilon(x,\theta_{0}),\quad S_{1}(x)=\varepsilon^{\prime}_{\theta}(x,\theta_{0}).$ (18)
Then, following the procedures discussed above, one can construct the weights $Q_{0}(x)$ and $Q_{1}(x)$ and calculate the quantities $Y_{i}$ according to Eq. (14) for the given signal $\varepsilon(x,\theta)$. The dependence of the linear estimator defined by Eq. (15) on the true value of the parameter $\theta$ is, in general, also nonlinear and biased, $\bar{\theta}=\Theta(\theta)\neq\theta$. Assuming that the dependence is monotonically increasing in certain vicinity of the reference value $\theta_{0}$, one can construct its inverse and build the estimator as
$\bar{\theta}_{\text{nonlin}}=\Theta^{-1}(\bar{\theta}).$ (19)
Practically, the procedure means plotting the calibration curve $\bar{\theta}=\Theta(\theta)$ and using it for finding the parameter value $\theta=\Theta^{-1}(\bar{\theta})$ for the value $\bar{\theta}$ calculated from the input data $\varepsilon(x,\theta)$ (Fig. 1(d)).
II.3 Combining sensing and SOFI
Practically, such approach would mean that one takes the binned signal density $f(\mathbf{r},t,x_{i})$, specified for a discrete set of values $x_{i}$, and performs analysis of frame-wise temporal statistics to each bin $x_{i}$ by considering the time series $\{f(\mathbf{r},t,x_{i}\}_{t}$. However, such time series is unlikely to satisfy the requirement of having enough photons in each frame, which is important for reliable application of SOFI.
Such issue can be resolved by the above-discussed “compression” of a full profile over $x$ into several scalar functionals. Following Eq. (14), one can define the signals, integrated over two spectral or temporal windows (indexed as $j=0,1$) within each frame, as follows:
[…]
where $Y_{kj}$ is calculated according to Eq. (14) for $\theta=\theta_{k}$.
@@ II.3 Combining sensing and SOFI
$C_{j}^{(n)}(\mathbf{r})=C_{t}^{(n)}[F_{j}(\mathbf{r},t)]\\ =\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{n}Y_{kj}^{n}C_{t}^{(n)}[s_{k}(t)].$ (22)
The quantities $C_{0}^{(n)}(\mathbf{r})$ and $C_{1}^{(n)}(\mathbf{r})$ have the meaning close to $Y_{0}$ and $Y_{1}$ and can be used for estimation of the local parameter value $\theta(\mathbf{r})$ similarly to Eq. (15). If the deviations $|\theta_{k}-\theta_{0}|$ of the parameter values $\theta_{k}$ at the emitters’ positions from the reference value $\theta_{0}$ are small, the factor $Y_{kj}^{n}$ in Eq. (22) can be decomposed in the way similar to Eq. (17):
$Y_{kj}^{n}\approx A_{j0}^{n}+nA_{j0}^{n-1}A_{j1}(\theta_{k}-\theta_{0}).$ (23)
Comparison of Eqs. (17) and (23) suggests that one can construct the estimator for the local value $\theta(\mathbf{r})$ of the environment parameter as
$\bar{\theta}^{(n)}(\mathbf{r})=\theta_{0}\\ +\frac{C_{1}^{(n)}(\mathbf{r})A_{00}^{n}-C_{0}^{(n)}(\mathbf{r})A_{10}^{n}}{n[C_{0}^{(n)}(\mathbf{r})A_{10}^{n-1}A_{11}-C_{1}^{(n)}(\mathbf{r})A_{00}^{n-1}A_{01}]}.$ (24)
Substituting Eqs. (22) and (23) into Eq. (24), one can see that the constructed estimator indeed represents the parameter distribution, blurred by the diffraction:
$\bar{\theta}^{(n)}(\mathbf{r})=\theta_{0}\\ +\frac{\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{n}C_{t}^{(n)}[s_{k}(t)](\theta_{k}-\theta_{0})}{\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{n}C_{t}^{(n)}[s_{k}(t)]}.$ (25)
The effect of blurring reduction for higher-order cumulants, characteristic for SOFI, is still present in Eq. (25): the PSF $U$ is raised to the power $n$.
For larger deviations $\theta-\theta_{0}$, the nonlinearity of the response should be taken into account similarly to Eq. (19). For a uniform distribution of the parameter $\theta$ ($\theta(\mathbf{r})=const=\theta$), one has $C_{j}^{(n)}(\mathbf{r})\propto Y_{j}^{n}$ and can find the (nonlinear) dependence $\bar{\theta}^{(n)}=\Theta_{n}(\theta)$ of the estimator $\bar{\theta}^{(n)}$ on the parameter value $\theta$, calculated according to Eq. (24) with the substitution $C_{j}^{(n)}(\mathbf{r})\mapsto Y_{j}^{n}$. Then, in the general case of non-uniform parameter distribution, the nonlinear estimator will be
$\bar{\theta}^{(n)}_{\text{nonlin}}(\mathbf{r})=\Theta_{n}^{-1}(\bar{\theta}^{(n)}(\mathbf{r})),$ (26)
where $\bar{\theta}^{(n)}(\mathbf{r})$ is calculated according to Eq. (24). General properties of the estimator are discussed in Appendix C.

### v2

@@ Abstract
We propose a method of spatial resolution enhancement in metrology (thermometry, magnetometry, pH estimation, and similar methods) with blinking fluorescent nanosensors by combining sensing with super-resolution optical fluctuation imaging (SOFI). To demonstrate efficiency of this approach, both numerical simulations and a model experiment with laser diodes, which model fluctuating nanoemitters, and intentional blurring of the image are performed. The 2nd, 3rd, and 4th order cumulant images provide improvement of the contrast and enable successful reconstruction of smaller features of the modeled temperature (or any other physical parameter) distribution relatively to the intensity-based approach. We believe that blinking fluorescent sensing agents being complemented with the developed image analysis technique could be utilized routinely in the life science sector for recognizing the local changes in the spectral response of blinking fluorophores, e.g. delivered targetly to the wanted cell or even organelle. It is extremely useful for the local measurements of living cells’ physical parameters changes due to applying any external “forces”, including disease effect, aging, healing, or response to the treatment.
@@ I Introduction
Fluorescent nanosensors represent a powerful tool for measurement of temperature, magnetic field, pH, and other parameters with high accuracy at small scales [1, 2, 3, 4, 5, 6]. In particular, the dependence of their fluorescent signal on the mentioned parameters can be utilized for investigation of processes in living cells [2, 7, 8]. To ensure high spatial resolution, one needs to localize the nanoemitters precisely and to distinguish signals from closely located nanosensors. This requirement can be fulfilled, for example, by placement of fluorescent nanodiomonds or quantum dots at a manipulator needle [9, 10, 11, 12, 13]. This approach requires precise mechanical control and does not allow to analyze environment parameters at several points of the sample simultaneously. For a large number of nanosensors, simultaneously attached to the sample, one can apply deterministic super-resolution approaches based on selective (localized) excitation of emitters (stimulated emission depletion (STED), ground state depletion (GSD), reversible saturable optical fluorescence transitions between spin levels (spin-RESOLFT), structured illumination imaging) [14, 15, 16, 17, 18], as well as stochastic methods (stochastic optical reconstruction microscopy (STORM), photoactivated localization microscopy (PALM), points accumulation for imaging in nanoscale topography (PAINT)) [19, 20, 21, 22], ensuring independent localization of single emitters due to their blinking. Those approaches are also quite sophisticated and require the nanosensors to remain in the inactive state most of the time, thus increasing needed exposition time and potential photo-toxicity.
@@
Combining SOFI with sensing is not as trivial as it may seem at the first glance. First, we focus on the case when the parameter of interest may vary at the scales smaller that the classical resolution limit, determined by the point spread function (PSF) of the used microscope. Therefore, the assumptions of slowly varying parameters, used for example in balanced SOFI (bSOFI) [29], are not generally valid here. Second, SOFI imposes less restrictions on separate frames of the acquired timeseries in comparison with single-molecule localization techniques [14, 15, 16, 17]: it allows smaller numbers of detected photons per frame and higher probability of closely positioned emitters being active during the same frame. Thus, one cannot reliably perform reconstruction of the sensed environmental parameter for a single-frame dataset or for a separate localized emitter as it is done in spectrally resolved STORM, PAINT, and PALM techniques [30, 31, 32, 33, 34]. Third, SOFI requires analysis of the whole acquired timeseries, relies on specific temporal fluctuations of the signal, and yields nonlinear dependence of the response on the emitters’ brightness [29, 35, 36]. Therefore, conversion of each separate frame into the investigated parameter map before processing of the entire timeseries would alter the fluctuations and preclude further application of SOFI for spatial resolution enhancement. On the other hand, application of SOFI to multichannel frame data (for example, spectra) would tremendously enhance the shot noise, while nonlinearity of the response would require additional treatment.
To resolve the outlined issues and combine SOFI with sensing, we propose to “compress” each frame to a two-channel image, apply SOFI to the timeseries corresponding to each of the two channels separately, and infer the sensed parameter map from the obtained cumulant images for those channels. Linearity of the compression ensures preservation of the temporal fluctuations of the signal and applicability of SOFI to spatial resolution enhancement. The idea of using two channels for single-parameter inference is closely related to ratiometric imaging [37, 38], where two spectral channels can be used for self-referencing [39] to reduce the adverse effect of absolute intensity fluctuations [37, 39, 40]. Nonlinearity of the resulting response of SOFI [29, 35, 36] can be compensated analytically by inversion of the nonlinear relation between the sensed parameter of interest and cumulant values.
@@
The idea of combined usage of blinking statistics of fluorophores for super-resolution and certain other degree of freedom for getting additional information is somewhat close to so called “color SOFI” [41, 42, 43, 44, 45] — super-resolution technique, which applies SOFI to several (typically two) spectral channels. However, the tasks, solved by those two approaches, are different. Color SOFI deals with discrimination between several (two or three) discrete types of fluorescent emitters, serving as biological labels. In contrast, the method proposed in this paper, deals with continuous-variable sensing: the emitters are of the same type, but their fluorescent signal differs due to variations of local environment.
@@
To demonstrate principles of the proposed approach and show its feasibility, we first applied in to multiple synthetic (simulated) datasets and quantified the performance and resolution enhancement. Information from both auto- and cross-cumulants [24] with zero time lag was combined to reduce the influence of the shot noise and improve the accuracy of sensing. The mean square error of the sensed parameter estimation scaled as $1:0.28:0.0075$ for the 1st, 2nd, and 4th order cumulants for the specific considered model. To get closer to real-world testing of the approach in non-ideal conditions (without exact prior knowledge of PSF and with measurement imperfections), we also performed a model experiment mimicking blinking fluorescent nanoemitters. We used classical photodiods with the controlled signal imitating random switching between “bright” and “dark” states. Similarly to model (proof-of-principle) super-resolution experiments in quantum imaging [46, 47, 48, 49, 50], the image was intentionally blurred (in our case, by inaccurate focusing on the detection plane) so that the spots produced by neighbor “emitters” were strongly overlapping. Model “temperature” of the “emitters” was encoded in the “fluorescence lifetime” — the delay between the “excitation pulse” and the emission response. The subsequent application of the proposed SOFI-based sensing approach allowed to enhance spatial resolution approximately twice even for the simplest version of SOFI with zero time lag autocumulants.
@@
The paper is organized as follows. In Section II, we discuss basic ideas of the approach and list the procedures constituting the proposed technique of spatially super-resolving sensing. Section III is devoted to application of the approach to synthetic datasets and quantification of the resolution (accuracy of parameter estimation) improvement. Section IV describes the model experiment, data acquisition, and reconstruction results for the model “temperature” distribution. Conclusions on the results and possible directions of further research are presented in Section IV.5.
@@ II.1 Super-resolution optical fluctuations imaging
Further, one can define also cross-cumulants and introduce non-zero time lags [24]:
$C_{t}^{(n)}[F(\mathbf{r}^{(1)},t+\tau_{1}),\ldots,F(\mathbf{r}^{(n)},t+\tau_{n})]\\ \propto\sum_{k}U^{n}(\bar{\mathbf{r}}-\mathbf{r}_{k})\varepsilon_{k}^{n}C_{t}^{(n)}[s_{k}(t+\tau_{1}),\ldots,s_{k}(t+\tau_{n})],$ (10)
where $\bar{\mathbf{r}}=\sum_{i=1}^{n}\mathbf{r}^{(i)}/n$.
To increase spatial resolution, SOFI deals with stochastic fluctuations of the total fluorescence signal, integrated over each frame duration and certain spectral range, rather than on spectral or temporal intensity distribution inside a frame. The technique itself does not impose any specific constraints on time or spectral windows, used during signal acquisition within frames. That property is crucial for the proposed sensing approach and was efficiently used in the color SOFI [41, 44]. It is worth noting, that there also exists an alternative approach of examining the emitters’ surrounding by analysis of the parameters of fluctuation already used for SOFI (for example, on-time ratio or blinking frequency) [29, 51, 52] rather than additional “degrees of freedom”.
@@ II.2 Sensing with fluorescent nanoparticles
Sensing capabilities of fluorescent nanoparticles can be provided by dependence of their fluorescence lifetime, spectrum, or total intensity on the local parameters of the environment: temperature, pH, pressure, etc. [1, 2, 53, 54, 55, 56, 57, 58, 59, 60, 61] (Fig. 1). For example, carbon nanodots possess spectral dependence of fluorescence on pH and temperature [60] and can exhibit blinking required for SOfI [45]. Surface-induced charge-state switching of color centers in nanodiamonds [62, 63, 64, 65, 66, 67] enables combination of their vast sensing capabilities [1, 2, 53, 54, 55, 56, 57] with SOFI.
In other words, variation of the parameter of interest can change the total intensity of the fluorescent signal and (or) its distribution over certain parameter (variable) $x$ (wavelength, emission delay, etc.) within the frame. Variations of molecular surrounding [37, 39, 40] and blinking of the emitters (required for SOFI applicability) [29] introduce significant noise to the total (absolute) intensity of fluorescence. For that reason, further we focus on the latter case when the sensed environmental parameter influences the distribution of the fluorescent signal. Let $f(\mathbf{r},t,x)$ be the density of the signal over the variable $x$, connected with the total fluorescent signal $F(\mathbf{r},t)$ for the frame $t$ by the relation
@@
Figure 1: Sensing an environment parameter $\theta$ by its influence on fluorescence signal. The value of the parameter can be inferred from changes of the spectrum (a), fluorescence lifetime (b), or total intensity (c). The latter case is likely to yield worse sensing accuracy because of blinking-induced noise of absolute intensity [29]. The full spectral profile (a) can be compressed into two channels: with the optimal weights described in Appendix A (d) and with practical spectral windows having Gaussian profiles of sensitivity $Q_{1}(x)$ and $Q_{2}(x)$ (e). The channel $Y_{0}$ is defined as $Y_{1}+Y_{2}$ and corresponds to the weight $Q_{0}(x)=Q_{1}(x)+Q_{2}(x)$. The full temporal profile (b) can be compressed into two channels by using a simple threshold filter (f).
@@
where the dependence on the $k$-th emitter intensity $\varepsilon_{k}(x,\theta)$ on the variable $x$ (wavelength or delay of the fluorescence signal) is affected by the local value of the environment parameter $\theta=\theta_{k}$ in the vicinity of the $k$-th emitter. To infer the local values $\theta(\mathbf{r})$ of the parameter in question, one needs to specify certain model for the dependence of $\varepsilon_{k}(x,\theta)$ on $\theta$. Typically, the emitters are assumed to be described by the same model $\varepsilon_{k}(x,\theta_{k})=\varepsilon_{k}\varepsilon(x,\theta_{k})$ up to a constant emitter-dependent brightness factor $\varepsilon_{k}$.
When trying to combine SOFI with sensing, it is tempting to apply cumulants construction to each value of $x$ (spectral or temporal channel) independently. However, as we discuss in the next section, splitting the fluorescent signal into a large number of channels leads to low photon numbers, significant shot noise, and inefficiency of high-order cumulants usage in SOFI. On the other hand, reconstruction of the whole distribution $\varepsilon(x,\theta(\mathbf{r}))$ is not needed to infer the spatial map of a single parameter $\theta(\mathbf{r})$. As we show here, the whole intensity profile can be “compressed” to just two scalar values with their ratio being sufficient for estimation of the parameter $\theta$. Accumulation of the intensity into two such channels effectively reduces the shot noise and restores applicability of SOFI for spatial super-resolution.
For now, let us consider sensing by a single probe nanoparticle localized at the position $\mathbf{r}_{0}$ and subjected to the environment with the parameter value $\theta$:
$f(\mathbf{r},t,x)=U(\mathbf{r}-\mathbf{r}_{0})\varepsilon_{0}(x,\theta)s_{0}(t).$ (13)
To eliminate additional inaccuracies, caused by blinking of the emitter, finite data acquisition time, or unknown brightness $\varepsilon_{0}$, one can work with normalized signal by dividing its density over the total value:
[…]
The idea of analyzing ratios instead of absolute values of the signal for reduction of fluctuations effects is widely used in ratiometric imaging [37, 38]. One can notice that the normalized signal depends on the normalized density of the emitter’s signal and is determined by the parameter of interest $\theta$, but does not include irrelevant geometrical information or the emitter brightness.
Instead of considering the whole normalized profile $\tilde{f}(\mathbf{r}_{0},t,x)$ for a specific frame $t$ and the emitter position $\mathbf{r}_{0}$, one can construct a quantity
[…]
for certain weight functions $Q_{0}(x)$ and $Q_{1}(x)$ and consider it as a “compressed” version of the measured signal (Fig. 1(d-f)). As shown in Appendix A, there exists the (locally) optimal choice of the weights (Fig. 1(d)), for which the quantity $Z$ contains exactly the same amount of information about the sensed parameter $\theta$ as the whole intensity profile $\tilde{f}(\mathbf{r}_{0},t,x)$ (i.e., represents sufficient statistics).
From practical point of view, Eq. (15) defines two detection channels $Y_{0}$ and $Y_{1}$ and proposes to use their ratio $Z$ for inference of $\theta$. In a real experiment, implementation of the optimal detection channels (with the optimal weight functions $Q_{0}(x)$ and $Q_{1}(x)$) can be complicated or impossible. Alternatively, one can just use the signals from two practically available channels $Y_{1}$ and $Y_{2}$ with their weight functions $Q_{1}(x)$ and $Q_{2}(x)$ (for example, corresponding to the signal accumulated over two spectral or temporal windows — Fig. 1(e,f)), define $Y_{0}=Y_{1}+Y_{2}$ (and $Q_{0}(x)=Q_{1}(x)+Q_{2}(x)$), and construct $Z$ according to Eq. (15). Strictly speaking, such measurement-induced choice of the weights is, generally, non-optimal and may lead to loss of information relatively to the full profile analysis.
Substituting Eq. (14) into Eq. (15), one can find the relation between the parameter $\theta$ and the compressed signal $Z$:
[…]
If the a priori constraints $\theta_{\text{min}}\leq\theta\leq\theta_{\text{max}}$ on the sensed $\theta$ ensure monotonicity of the dependence $\zeta(\theta)$, the function $\zeta(\theta)$ can be inverted and the estimator (map of the signal to the parameter of interest) can be constructed as
$\bar{\theta}=\zeta^{-1}(Z).$ (17)
For a particular case of linear dependence of the emitter’s signal density on the parameter of environment in the region of interest $[\theta_{\text{min}},\theta_{\text{max}}]$
$\varepsilon(x,\theta)=S_{0}(x)+S_{1}(x)\theta,$ (18)
the estimator can be constructed analytically:
$Z=\frac{A_{10}+A_{11}\theta}{A_{00}+A_{01}\theta}\quad\Rightarrow\quad\bar{\theta}=\frac{A_{00}Z-A_{10}}{A_{11}-A_{01}Z},$ (19)
where
[…]
It is worth mentioning that for the optimal weights $Q_{0}(x)$ and $Q_{1}(x)$, defined in Appendix A, we have $A_{ij}=\delta_{ij}$ and $\bar{\theta}=Z$.
II.3 Combining sensing and SOFI
Practically, such approach would mean that one takes the binned signal density $f(\mathbf{r},t,x_{i})$, specified for a discrete set of values $x_{i}$, and performs analysis of frame-wise temporal statistics for each bin $x_{i}$ by considering the time series $\{f(\mathbf{r},t,x_{i}\}_{t}$. However, such time series is unlikely to satisfy the requirement of having enough photons in each frame, which is important for reliable application of SOFI. (While there have already exist methods capable of resolving shot-noise induced systematic artifacts in SOFI [68], the statistical inaccuracy inevitably grows for higher-order cumulants).
Such issue can be resolved by the above-discussed “compression” of a full profile over $x$ into several scalar quantities. Following Eq. (15), one can define the signals, integrated over two spectral or temporal windows (indexed as $j=0,1$) within each frame, as follows:
@@ II.3 Combining sensing and SOFI
is the $j$-th channel signal of the $k$-th emitter subjected to the environmental parameter value $\theta=\theta_{k}\equiv\theta(\mathbf{r}_{k})$.
@@
$C_{j}^{(n)}(\mathbf{r})=C_{t}^{(n)}[F_{j}(\mathbf{r},t)]\\ =\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})\varepsilon_{k}^{n}Y_{j}^{n}(\theta_{k})C_{t}^{(n)}[s_{k}(t)].$ (24)
In the way, inherent to SOFI, the blurring is reduced by raising the PSF $U$ in the power $n$. The expressions for the quantities $C_{0}^{(n)}(\mathbf{r})$ and $C_{1}^{(n)}(\mathbf{r})$ differ only by the factors $Y_{0}^{n}(\theta_{k})$ and $Y_{1}^{n}(\theta_{k})$ in the sums, while the leading contribution to the sum is provided by the emitters with $\mathbf{r}_{k}\approx\mathbf{r}$. Therefore, their ratio can be considered as an approximation of the local value of the compressed signal $Z(\mathbf{r})$:
$Z(\mathbf{r})\equiv\zeta(\theta(\mathbf{r}))\approx Z_{n}(\mathbf{r})\equiv\left[\frac{C_{1}^{(n)}(\mathbf{r})}{C_{0}^{(n)}(\mathbf{r})}\right]^{1/n}.$ (25)
For even $n$, the information about the sign of $Z(\mathbf{r})$ can be lost, since $Y_{j}^{n}(\theta_{k})\geq 0$ for even $n$ regardless of $\operatorname{sign}Y_{j}(\theta_{k})$. However, that limitation is typically irrelevant in practice, because non-negative weights $Q_{j}(x)\geq 0$ for integration over spectral or temporal windows ensure $Y_{j}\geq 0$.
The resulting estimator for the $n$-th order SOFI combined with sensing is
$\bar{\theta}^{(n)}(\mathbf{r})=\zeta^{-1}\left(Z_{n}(\mathbf{r})\right).$ (26)
The estimator possesses the following properties justifying its application (see Appendix B):
  1. 1.
    For a uniform (constant) parameter distribution $\theta(\mathbf{r})=\theta_{0}$, the estimator reproduces the correct parameter value for all orders $n$:
    $\bar{\theta}^{(n)}(\mathbf{r})=\theta_{0}\text{ for all {\hbox{\mathbf{r}}} and {\hbox{n}}}.$ (27)
  2. 2.
    If all the emitters are subjected to the environment parameter $\theta$ limited to certain range: $\theta_{k}\in[\theta_{\text{min}},\theta_{\text{max}}]$ for all $k$, the estimator will also belong to that range:
    $\bar{\theta}^{(n)}(\mathbf{r})\in[\theta_{\text{min}},\theta_{\text{max}}]\text{ for all {\hbox{\mathbf{r}}} and {\hbox{n}}}.$ (28)
  3. 3.
    For small variations of the parameter $\theta$ (i.e., in the linear approximation), the estimator will reproduce a weighted combination of the parameter values $\theta_{k}$ for the emitters in the vicinity of the point of interest:
    $\bar{\theta}^{(n)}(\mathbf{r})\approx\frac{\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})C_{t}^{(n)}[s_{k}(t)]\varepsilon_{k}^{n}\theta_{k}}{\sum_{k}U^{n}(\mathbf{r}-\mathbf{r}_{k})C_{t}^{(n)}[s_{k}(t)]\varepsilon_{k}^{n}}.$ (29)
    For larger $n$, the estimate becomes more localized due to raising the PSF $U$ to the power $n$.